# DiVERT: Distractor Generation with Variational Errors Represented as Text for Math Multiple-choice Questions

### v1

@@ Abstract
Nigel Fernandez1*, Alexander Scarlatos1*, Simon Woodhead2, Andrew Lan1 1University of Massachusetts Amherst, 2Eedi {nigel,ajscarlatos,andrewlan}@cs.umass.edu, simon.woodhead@eedi.co.uk
@@ 1 Introduction
The typical approach to MCQ distractor development primarily relies on extensive human effort, which can be burdensome for educators, which motivated the development of automated approaches. Prior work on automated distractor generation primarily focuses on i) cloze tasks in language learning to assess vocabulary recall or grammatical knowledge and ii) reading comprehension question answering to assess comprehension of a given text or article Alhazmi et al. (2024). Approaches include using knowledge graphs, encoder-decoder models, and with help from large language models (LLMs). See Supplementary Material A for a more detailed review of related work.
@@
For other subjects where common tasks require (possibly complex) reasoning ability, such as math, there exist relatively few approaches to automated MCQ distractor generation. In these domains, generating high-quality distractors is more challenging than in language learning or reading comprehension: distractors need to reflect abstract mathematical misconceptions and/or procedural errors in the mathematical reasoning process. Earlier works either use symbolic, rule-based approaches to generate distractors Tomás and Leal (2013); Prakash et al. (2023), which have limited generalizability beyond template-based questions, or sample incorrect generations during a math problem solving process as distractors Dave et al. (2021). More recently, Feng et al. (2024); Scarlatos et al. (2024) investigate a wide array of LLM-based approaches for math MCQ generation, particularly focusing on the plausibility of distractors, i.e., how likely is a distractor going to be selected among real students.
@@ 1.1 Contributions
    We introduce DiVERT  [^1] We will make our code publicly available. (Distractor Generation with Variational Errors Represented as Text), a novel variational approach to jointly learn error representations in math MCQs as well as generate distractors corresponding to errors.
@@ 2 Problem Formulation
We select distractor generation as the downstream task since performance on this task is easily quantifiable; however, there are other meaningful tasks, such as feedback message and tutoring dialogue turn generation, where knowing what error a student made is key and significantly improves LLMs’ performance on these tasks Jurenka et al. (2024), which we leave for future work.
@@ 3.1 Variational Approach
We face two key technical challenges: First, errors can be represented in different formats: discrete, as a finite set of error categories, continuous, as latent vectors that represent a distribution of errors, or textual, with errors represented as natural language explanations. Considering use cases in real-world educational scenarios, such as providing feedback to students and teachers and informing a chatbot in tutoring sessions, we choose to represent errors as text: An error behind a distractor in math MCQs is characterized by a sequence of textual tokens, as shown in Figure 1. Second, for the same MCQ stem, there can often be many plausible errors among real students. For the example in Figure 1, the correct solution approach involves several mathematical reasoning steps, where at each step (e.g., performing fraction multiplication), there are multiple errors that real students can make (e.g., multiplying the numerator and denominator, not multiplying the denominator, etc.) The error generation model $g^{\text{err}}$ must be able to capture the space of all plausible errors, which is challenging.
To address these challenges, we use a variational approach for our task and jointly learn $g^{\text{err}}$ and $g^{\text{dis}}$. The advantage of such an approach is that during training, we can train on errors sampled from the error distribution, which gives us a better shot at capturing the (possibly large) error space. Shown in Figure 1, DiVERT has the following components:
  1. 1.
    The error prior model $p(e|s)$: This model generates a textual explanation of an error that students can make given the MCQ’s stem.
  2. 2.
    The controllable distractor generation model $p(d|s,e)$: This model generates a distractor that is the resulting incorrect answer after making a specific error in this question.
  3. 3.
    An error identifier model $q(e|s,d)$: This approximate posterior model identifies the error behind a question-distractor pair.
We note that there are can be many choices for the variational distribution $q(\cdot)$; we choose $q(e|s,d)$ since it can be useful in practice and can serve as a sanity check tool. We leave experimenting with other useful approximate distributions such as $q(e|t)$ or even simply $q(e)$ for future work.
3.2 DiVERT
As discussed above, to promote interpretability in error representations, we parameterize all distributions in our variational approach using LLMs. We use a different LLMs for each distribution; each is fine-tuned separately from the same base LLM using QLoRA Dettmers et al. (2024). We denote them as $\text{LLM}^{e}$, $\text{LLM}^{d}$, and $\text{LLM}^{q}$.
Approximate Inference. Since exact inference of the posterior distribution is computationally intractable, $q_{\phi}(e|s,d)$ is an approximation of the true posterior. We perform approximate inference by maximizing the evidence lower bound (ELBO) on the observed data log-likelihood, given by
$\log p_{\theta}(d|s)$ $=\mathcal{L}(\theta,\phi)\geq\text{ELBO}(d|s)$
@@ 3.2 DiVERT
where the model parameters $\theta$ consist of tunable parameters in the error prior, $\theta_{e}$, the controllable distractor generator $\theta_{d}$, and the variational error identifier model $\phi$. The parameter $\beta>0$ controls the balance between the distractor reconstruction loss and the KL divergence between the approximate posterior and the prior. Next, we detail a series of novel methods that we introduce to enable DiVERT to perform our two tasks.
@@
Differentiable Learning through Discrete Tokens. The ELBO in Equation 3.2 can easily be approximated by Monte Carlo simulation, with samples drawn from $q_{\phi}(e|s,d)$ (see Supplementary Material B). However, since we parameterize $q_{\phi}$ with an LLM, these samples are sequences of discrete text tokens. Therefore, we cannot simply use the reparameterization trick to sample from $q_{\phi}$. Using a similar approach to Liu et al. (2023), we use soft tokens, i.e., differentiable versions of hard, discrete text tokens during training to enable the flow of gradients. Specifically, for the $k$-th error token $\hat{e}_{k}$ generated by $\text{LLM}^{q}$, we replace its discrete embedding $\mathbf{v}_{\hat{e}_{k}}$ with $\sum_{j}p_{k,j}\mathbf{v}_{j}$, where $p_{k,j}$ is the probability of generating token $j$ at position $k$ from $\text{LLM}^{q}$ and $\mathbf{v}_{j}$ is the embedding vector of token $j$ in the vocabulary of $\text{LLM}^{q}$. This approximation enables us to sample discrete tokens while differentiating through continuous vectors during training. We calculate $p_{k,j}=\operatorname{softmax}(\mathbf{z}_{k}/\lambda)_{j}$, where $\mathbf{z}_{k}$ is the output logit vector from $\text{LLM}^{q}$ at position $k$ and $\lambda>0$ is a temperature parameter. Following prior work Jang et al. (2017), we initialize $\lambda$ to $1$ and exponentially anneal it to $0.1$ during training. This process means that the soft tokens are smooth when training starts to enable better gradient flow, while later becoming closer to hard tokens.
@@
Initialization with Supervised Fine-tuning. Despite LLMs exhibiting better and better mathematical reasoning capabilities, their capacity in understanding errors Sonkar and Baraniuk (2023) and generating distractors that correspond to errors Feng et al. (2024) remains surprisingly poor. Therefore, we solicit a collection of error labels behind question-distractor pairs, $(s,d)$, from math educators. Then, we fine-tune all three LLMs in their respective formats using this annotation data to initialize DiVERT’s three LLM components. Without this step, base LLMs, even ones that perform well on question answering tasks, generate low-quality error explanations that hurt performance, especially during early training stages. In Section 5, we demonstrate that using a small portion of error labels for warm-up fine-tuning helps DiVERT’s ability to learn from unlabeled $(s,d)$ pairs.
@@
Q Regularization. Since the space of plausible errors can be large, we add a regularization term in our loss function to prevent the approximate posterior distribution $q_{\phi}(e|s,d)$ from deviating too much compared to its initialized version after supervised fine-tuning, $q_{\phi_{\text{init}}}$. This term, weighted by a balancing parameter $\alpha>0$, controls the amount of exploration in our error tokens sampled from $q_{\phi}(e|s,d)$ and prevents it from diverging during training. Concretely, the regularization is given by
@@
with our final training objective becoming
$\mathcal{L}(\theta,\phi)+\alpha\mathcal{L}_{\text{reg}}.$ (3)
@@
Overgenerate-and-rank for Errors and Distractors. At test time, we first generate a set of $N_{e}$ errors, denoted as $\hat{E}$, through $\text{LLM}^{e}$, using diverse beam search Vijayakumar et al. (2018) for decoding, to promote diversity among generated errors. Then, for each generated error $\hat{e}\in\hat{E}$, we generate $N_{d}$ distractors through $\text{LLM}^{d}$ using standard beam search since the distractors exhibit much less variation under a specific error. Finally, we rank all $N_{e}\times N_{d}$ candidate distractors in $\hat{D}$ by their associated beam scores and select the top-$K$.
@@ 4.1 Dataset Details
We work with a real-world math MCQ dataset containing $1,434$ MCQs written in English, each with a set of $3$ distractors. We collect error labels from middle school math teachers for each question-distractor pair, explaining why a student may select that distractor. The questions are designed primarily to assess students aged between $10$ to $13$, on $41$ unique subtopics, including “Basic Arithmetic”, “Fractions”, and “Solving Equations”. We divide the dataset into train-val-test splits by questions to ensure no overlap in the MCQ stem across the splits, resulting in roughly a $72\%$-$16\%$-$12\%$ split over question-distractor pairs. See Supplementary Material D for statistics, and Supplementary Material G for MCQ examples.
@@ 4.2 Metrics
Error Evaluation. The open-ended and mathematical nature of errors makes automated text similarity metrics like ROUGE-L F1 Lin (2004) and BERTScore F1 Zhang et al. (2020) unsuitable. Therefore, we conduct a human evaluation of generated errors, which we detail in Section 6. For completeness, we report error evaluation on automated metrics in Supplementary Material E.
Distractor Evaluation. Following prior work on automated distractor generation Feng et al. (2024), we use alignment-based metrics to measure how well the $K$ generated distractors align with ground-truth human-authored ones. The first metric, Exact match ($h_{e}$), measures whether all three human-authored distractors in $D$ are matched exactly by some subset of the generated distractors $\hat{D}$. Similarly, Partial match ($h_{p}$) measures whether at least one generated distractor matches human-authored ones. Concretely, these binary metrics are defined as $h_{e}(D,\hat{D})=1$ if $D\subseteq\hat{D}$, and $h_{p}(D,\hat{D})=1$ if $\hat{D}\cap D\neq\emptyset$. The third, continuous metric, Proportional match ($h_{n}$), measures the portion of human-authored distractors that match generated ones, defined as $h_{n}(D,\hat{D})=|\hat{D}\cap D|/3$. We compute all metrics averaged across all MCQs in the test set and report percentages. We vary $K\in\{3,10\}$, similar to the setup for metrics such as MAP@K. The Proportional match metric is most important since it is more robust than the other two.
@@ 4.3 Baselines
GPT-4o Zero-shot CoT Feng et al. (2024) $2.77$ $68.75$ $33.33$ $20.83$ $81.25$ $50.46$
GPT-4o kNN 25.69 83.33 55.09 $36.80$ $88.88$ $63.88$
@@
DisSearch-D $13.88$ $75$ $41.66$ $36.11$ $90.27$ $64.12$
DisSearch-ED CoT $13.88$ $78.47$ $43.51$ $37.5$ $\underline{90.97}$ $\underline{66.66}$
DisSearch-ED CoT Pipeline $\underline{15.27}$ $75$ $43.28$ $\underline{38.19}$ $88.88$ $64.58$
DiVERT (ours) $13.19$ $\underline{81.25}$ $\underline{46.06}$ 42.36 91.66 68.75
Table 1: Performance on distractor generation for all approaches across all metrics. DiVERT, using an open-source base LLM with $7$B parameters, outperforms all baselines and performs on par with or better than the much larger and proprietary GPT-4o. Best performance is in bold and second best is underlined.
Prompting Baselines. We compare DiVERT to two prompting-based approaches proposed in Feng et al. (2024) using the state-of-the-art LLM GPT-4o OpenAI (2024). In contrast to our approach, these approaches generate all distractors in one pass, without considering errors behind them. The first is kNN, their best-performing approach, where we use the 3 most similar MCQs in the training set as in-context examples and teacher-written feedback as chain-of-thought reasoning to aid distractor generation. The second is CoT, where no in-context examples are given but the model is prompted to generate chain-of-thought Wei et al. (2022) reasoning before each distractor. In our CoT implementation, we prompt the model to generate textual error explanations as chain-of-thought so the errors can also be directly compared with those from our approach. We make other small updates to the prompts to include question tags and to generate either 3 or 10 distractors per question. All our prompts are shown in Supplementary Material H.
Fine-Tuning Baselines. We introduce three strong fine-tuning baseline approaches for distractor generation. The first is DisSearch-D, where we fine-tune the base LLM to directly generate a distractor from the question stem, i.e., training $p(d|s)$ on all question-distractor pairs. The second is DisSearch-ED CoT, where we fine-tune the base LLM to generate the error first, as chain-of-thought Wei et al. (2022) reasoning, followed by the distractor, i.e., training $p(e,d|s)$ on all question-distractor pairs. At inference time, we also use beam search to match the overgenerate-and-rank setup used in DiVERT. The third is DisSearch-ED CoT Pipeline, where we use the fine-tuned $p(d|s,e)$ and $p(e|s)$ models without variational training and test with the same two-step pipeline as DiVERT.

### v2

@@ Abstract
Nigel Fernandez1*, Alexander Scarlatos1*, Wanyong Feng1 Simon Woodhead2, Andrew Lan1 1University of Massachusetts Amherst, 2Eedi {nigel,ajscarlatos,wanyongfeng,andrewlan}@cs.umass.edu simon.woodhead@eedi.co.uk
@@ 1 Introduction
The typical approach to MCQ distractor development primarily relies on extensive human effort, which can be burdensome for educators, which motivated the development of automated approaches. Prior work on automated distractor generation primarily focuses on i) cloze tasks in language learning to assess vocabulary recall or grammatical knowledge and ii) reading comprehension question answering to assess comprehension of a given text or article Alhazmi et al. (2024). Approaches include using knowledge graphs, encoder-decoder models, and with help from large language models (LLMs). See Section 7 for a more detailed review of related work.
@@
For other subjects where common tasks require (possibly complex) reasoning ability, such as math, there exist relatively few approaches to automated MCQ distractor generation. In these domains, generating high-quality distractors is more challenging than in language learning or reading comprehension: distractors need to reflect abstract mathematical misconceptions and/or procedural errors in the mathematical reasoning process. Earlier works either use symbolic, rule-based approaches to generate distractors Tomás and Leal (2013); Prakash et al. (2023), which have limited generalizability beyond template-based questions, or sample incorrect generations during a math problem solving process as distractors Dave et al. (2021). More recently, Feng et al. (2024); Scarlatos et al. (2024a) investigate a wide array of LLM-based approaches for math MCQ generation, particularly focusing on the plausibility of distractors, i.e., how likely is a distractor going to be selected among real students.
@@ 1.1 Contributions
    We introduce DiVERT 111 Code: https://github.com/umass-ml4ed/divert (Distractor Generation with Variational Errors Represented as Text), a novel variational approach to jointly learn error representations in math MCQs as well as generate distractors corresponding to errors.
@@ 2 Problem Formulation
We select distractor generation as the downstream task since performance on this task is easily quantifiable; however, there are other meaningful tasks, such as feedback message and tutoring dialogue turn generation, where current methods struggle to implicitly identify student errors McNichols et al. (2024); Scarlatos et al. (2024b), yet knowing what error a student made significantly improves LLMs’ performance on these tasks Jurenka et al. (2024). We leave exploring such tasks for future work.
@@ 3.1 Variational Approach
There can be numerous plausible errors among real students for each math MCQ, all of which may lead to different distractors. For example, in the MCQ in Figure 1 on fraction division, the distractor $\frac{7}{18}$ may result from a student exhibiting the error “multiplies numerators and denominators”, while the distractor $\frac{8}{9}$ may result from the error “adds numerators and denominators”. Without explicit error annotations, an error $e$ is a latent variable underlying a distractor $d$. Thus, the observed likelihood of a distractor given a question stem $s$ is
$p(d|s)=\textstyle\sum_{e\in\mathcal{E}}p(e|s)p(d|s,e),$ (1)
where $p(e|s)$ is the probability that the error $e$ can be made on question $s$, $p(d|s,e)$ is the probability that $d$ corresponds to the error $e$, and $\mathcal{E}$ is the space of all possible errors, which can be computationally intractable depending on how the errors are represented. While there are many potential representations of the errors, such as discrete categories or continuous latent vectors, we choose to represent each error as a sequence of textual tokens, so that the learned underlying representations of distractors are interpretable by both humans and language models.
Given this perspective, we naturally use a variational approach to learn an approximation of the large latent error space. Similar to variational auto-encoders Kingma and Welling (2014), we jointly learn the latent distribution and maximize the likelihood of the data by maximizing the evidence lower bound (ELBO) on the observed data log-likelihood, given by
$\log p_{\theta}$ $(d|s)\geq\text{ELBO}(d|s)=\mathcal{L}(\theta,\phi)$
@@
where $\theta_{e}$, $\theta_{d}$, and $\phi$ are the parameters of the prior model, distractor likelihood model, and variational model, respectively. The parameter $\beta>0$ controls the balance between the distractor reconstruction loss and the KL divergence between the approximate posterior and the prior.
To maximize the objective above, we jointly train three models, each parameterized by an LLM:
  1. 1.
    The error prior model $p_{\theta_{e}}(e|s)$: This model generates a textual explanation of an error that students can make given the MCQ’s stem.
  2. 2.
    The controllable distractor generation model $p_{\theta_{d}}(d|s,e)$: This model generates a distractor that is the resulting incorrect answer after making a specific error in this question.
  3. 3.
    An error identifier model $q_{\phi}(e|s,d)$: This approximate posterior model generates the error behind a question-distractor pair, and acts as the variational distribution during training.
We note that there can be many choices for the variational distribution $q(\cdot)$; we choose $q(e|s,d)$ since it can be useful in practice by allowing us to recover textual errors from distractors. We leave experimenting with other useful approximate distributions such as $q(e|t)$ or even simply $q(e)$ for future work. Next, we detail how we train all three models in a connected fashion, with the pipeline shown in Figure 1.
3.2 DiVERT
As discussed above, to promote interpretability in error representations, we parameterize all distributions in our variational approach using LLMs. We use a different LLM for each distribution; each is fine-tuned from the same base LLM using QLoRA Dettmers et al. (2024). We denote them as $\text{LLM}^{e}$, $\text{LLM}^{d}$, and $\text{LLM}^{q}$. Next, we detail a series of novel methods that we introduce to enable DiVERT to effectively perform the tasks of error learning and distractor generation.
Differentiable Learning through Discrete Tokens.
@@ 3.2 DiVERT
The ELBO in Equation 3.1 can easily be approximated by Monte Carlo simulation, with samples drawn from $q_{\phi}(e|s,d)$ (see Supplementary Material A). However, since we parameterize $q_{\phi}$ with an LLM, these samples are sequences of discrete text tokens. Therefore, we cannot simply use the reparameterization trick to sample from $q_{\phi}$. Using a similar approach to Liu et al. (2023), we use soft tokens, i.e., differentiable versions of hard, discrete text tokens during training to enable the flow of gradients. Specifically, for the $k$-th error token $\hat{e}_{k}$ generated by $\text{LLM}^{q}$, we replace its discrete embedding $\mathbf{v}_{\hat{e}_{k}}$ with $\sum_{j}p_{k,j}\mathbf{v}_{j}$, where $p_{k,j}$ is the probability of generating token $j$ at position $k$ from $\text{LLM}^{q}$ and $\mathbf{v}_{j}$ is the embedding vector of token $j$ in the vocabulary of $\text{LLM}^{q}$. This approximation enables us to sample discrete tokens while differentiating through continuous vectors during training. We calculate $p_{k,j}=\operatorname{softmax}(\mathbf{z}_{k}/\lambda)_{j}$, where $\mathbf{z}_{k}$ is the output logit vector from $\text{LLM}^{q}$ at position $k$ and $\lambda>0$ is a temperature parameter. Following prior work Jang et al. (2017), we initialize $\lambda$ to $1$ and exponentially anneal it to $0.1$ during training. This process means that the soft tokens are smooth when training starts to enable better gradient flow, while later becoming closer to hard tokens.
Initialization with Supervised Fine-tuning.
@@
Despite LLMs exhibiting better and better mathematical reasoning capabilities, their capacity in understanding errors Sonkar and Baraniuk (2023) and generating distractors that correspond to errors Feng et al. (2024) remains surprisingly poor. Therefore, we solicit a collection of error labels behind question-distractor pairs, $(s,d)$, from math educators. Then, we fine-tune all three LLMs in their respective formats using this annotation data to initialize DiVERT’s three LLM components. Without this step, base LLMs, even ones that perform well on question answering tasks, generate low-quality error explanations that hurt performance, especially during early training stages. In Section 5, we demonstrate that using a small portion of error labels for warm-up fine-tuning helps DiVERT’s ability to learn from unlabeled $(s,d)$ pairs.
Q Regularization.
@@
Since the space of plausible errors can be large, we add a regularization term in our loss function to prevent the approximate posterior distribution $q_{\phi}(e|s,d)$ from deviating too much compared to its initialized version after supervised fine-tuning, $q_{\phi_{\text{init}}}$. This term, weighted by a balancing parameter $\alpha>0$, controls the amount of exploration in our error tokens sampled from $q_{\phi}(e|s,d)$ and prevents it from diverging during training. Concretely, the regularization is given by
@@
with our final training objective to minimize becoming
$-\mathcal{L}(\theta,\phi)+\alpha\mathcal{L}_{\text{reg}}.$ (4)
@@
Overgenerate-and-rank for Errors and Distractors.
At test time, we first generate a set of $N_{e}$ errors, denoted as $\hat{E}$, through $\text{LLM}^{e}$, using diverse beam search Vijayakumar et al. (2018) for decoding, to promote diversity among generated errors. Then, for each generated error $\hat{e}\in\hat{E}$, we generate $N_{d}$ distractors through $\text{LLM}^{d}$ using standard beam search since the distractors exhibit much less variation under a specific error. Finally, we rank all $N_{e}\times N_{d}$ candidate distractors in $\hat{D}$ by their associated beam scores and select the top-$K$ Ashok Kumar et al. (2023).
@@ 4.1 Dataset Details
We work with a real-world math MCQ dataset from the Eedi learning platform [^2] https://eedi.com/us, containing $1,434$ MCQs written in English, each with a set of $3$ distractors. We collect error labels from middle school math teachers for each question-distractor pair, explaining why a student may select that distractor. The questions are designed primarily to assess students aged between $10$ to $13$, on $41$ unique subtopics, including “Basic Arithmetic”, “Fractions”, and “Solving Equations”. We divide the dataset into train-val-test splits by questions to ensure no overlap in the MCQ stem across the splits, resulting in roughly a $72\%$-$16\%$-$12\%$ split over question-distractor pairs. See Supplementary Material D for statistics, and Supplementary Material G for MCQ examples.
@@ 4.2 Metrics
Error Evaluation.
The open-ended and mathematical nature of errors makes automated text similarity metrics like ROUGE-L F1 Lin (2004) and BERTScore F1 Zhang et al. (2020) unsuitable. Therefore, we conduct a human evaluation of generated errors, which we detail in Section 6. For completeness, we report error evaluation on automated metrics in Supplementary Material E.
Distractor Evaluation.
Following prior work on automated distractor generation Feng et al. (2024), we use alignment-based metrics to measure how well the $K$ generated distractors align with ground-truth human-authored ones. The first metric, Exact match ($h_{e}$), measures whether all three human-authored distractors in $D$ are matched exactly by some subset of the generated distractors $\hat{D}$. Similarly, Partial match ($h_{p}$) measures whether at least one generated distractor matches human-authored ones. Concretely, these binary metrics are defined as $h_{e}(D,\hat{D})=1$ if $D\subseteq\hat{D}$, and 00 otherwise, while $h_{p}(D,\hat{D})=1$ if $\hat{D}\cap D\neq\emptyset$, and 00 otherwise. The third, continuous metric, Proportional match ($h_{n}$), measures the portion of human-authored distractors that match generated ones, defined as $h_{n}(D,\hat{D})=|\hat{D}\cap D|/3$. We compute all metrics averaged across all MCQs in the test set and report percentages. We vary $K\in\{3,10\}$, similar to the setup for metrics such as MAP@K. The Proportional match metric is most important since it is more robust than the other two.
@@ 4.3 Baselines
GPT-4o Zero-shot CoT Feng et al. (2024) $6.22_{\pm 2.18}$ $69.14_{\pm 3.97}$ $35.27_{\pm 1.74}$ $19.47_{\pm 3.30}$ $78.66_{\pm 3.71}$ $50.00_{\pm 1.96}$
GPT-4o kNN Feng et al. (2024) $\textbf{21.28}_{\pm 3.43}$ $\textbf{78.42}_{\pm 5.81}$ $\textbf{49.63}_{\pm 3.88}$ $33.47_{\pm 3.48}$ $85.14_{\pm 4.77}$ $60.19_{\pm 3.87}$
@@
DisSearch-D $13.74_{\pm 2.86}$ $74.13_{\pm 4.52}$ $41.76_{\pm 0.50}$ $34.12_{\pm 2.33}$ $86.17_{\pm 4.91}$ $61.41_{\pm 2.68}$
DisSearch-ED CoT $\underline{14.11}_{\pm 1.22}$ $73.18_{\pm 3.81}$ $42.14_{\pm 1.52}$ $\underline{36.21}_{\pm 1.21}$ $\underline{86.77}_{\pm 3.80}$ $\underline{62.83}_{\pm 2.52}$
DisSearch-ED CoT Pipeline $13.53_{\pm 1.74}$ $73.63_{\pm 4.05}$ $41.26_{\pm 1.87}$ $32.97_{\pm 3.48}$ $86.23_{\pm 3.88}$ $60.42_{\pm 2.80}$
DiVERT (ours) $13.37_{\pm 1.70}$ $\underline{76.33}_{\pm 4.33}$ $\underline{42.87}_{\pm 2.45}$ $\textbf{37.00}_{\pm 3.29}$ $\textbf{87.26}_{\pm 4.29}$ $\textbf{63.24}_{\pm 3.37}$
Table 1: Cross-validation performance on distractor generation for all approaches across all metrics. DiVERT, using an open-source base LLM with $7$B parameters, outperforms all baselines and performs on par with or better than the much larger and proprietary GPT-4o. Best performance is in bold and second best is underlined.
Prompting Baselines.
We compare DiVERT to two prompting-based approaches proposed in Feng et al. (2024) using the state-of-the-art LLM GPT-4o OpenAI (2024a). In contrast to our approach, these approaches generate all distractors in one pass, without considering errors behind them. The first is kNN, their best-performing approach, where we use the 3 most similar MCQs in the training set as in-context examples and teacher-written errors as chain-of-thought reasoning to aid distractor generation. We only use questions that have errors for all distractors as examples. The second is CoT, where no in-context examples are given but the model is prompted to generate errors as chain-of-thought Wei et al. (2022) reasoning before each distractor. We note that Feng et al. (2024) use teacher-written feedback instead of errors as CoT, but we use errors for a fairer comparison to our method. We make other small updates to the prompts to include question tags and to generate either 3 or 10 distractors per question. All our prompts are shown in Supplementary Material H. In Supplementary Material B, we show results from other baselines in Feng et al. (2024) on our data. We also perform a preliminary, small-scale comparison with the recent OpenAI o1 model OpenAI (2024b), which reportedly performs better than GPT-4o on mathematical reasoning tasks.
Fine-Tuning Baselines.
We introduce three strong fine-tuning baselines for distractor generation. The first is DisSearch-D, where we fine-tune the base LLM to directly generate a distractor from the question stem, i.e., training $p(d|s)$ on all question-distractor pairs. The second is DisSearch-ED CoT, where we fine-tune the base LLM to generate the error first, as chain-of-thought Wei et al. (2022) reasoning, followed by the distractor, i.e., training $p(e,d|s)$ on all question-distractor pairs. At inference time, we also use beam search to match the overgenerate-and-rank setup used in DiVERT. The third is DisSearch-ED CoT Pipeline, where we use the fine-tuned $p(d|s,e)$ and $p(e|s)$ models without variational training and test with the same two-step pipeline as DiVERT.